OpenStudy (anonymous):

can someone simplify log2^12 - log2^6 ???

OpenStudy (asadkarim7):

apply quotient law log 2^12/6 log 2^2= 1

OpenStudy (anonymous):

yes the difference of two logs with the same base can be regrouped into a single log of a quotient

OpenStudy (anonymous):

\[\log(2^{12})-\log(2^6) \rightarrow \log \frac{2^{12}}{2^6}\]

OpenStudy (anonymous):

well the answer is 1

OpenStudy (anonymous):

because u arrive at log2^2, and the inverse property loga^a^x suggests that the answer is 1

OpenStudy (anonymous):

I confused \[\log(2^{12})-\log(2^6) <> 1\] it equals 1.806!

OpenStudy (anonymous):

the answer should be \[\log(2^6)=1.806\]